The armature of a small generator consists of a flat, square coil with 160 turns and sides with a length of 1.55 cm . The coil rotates in a magnetic field of 7.60×10−2 T .

What is the angular speed of the coil if the maximum emf produced is 2.10×10−2 V ? (in rad/s)
Answer #1

Magnetic Flux through square coil, = BA cos wt
A = area of cross-section = 0.0155*0.0155 m2
w = angular speed, t =time
instantaneous emf = e(t) = - N d()/dt = - Nw (BA) [- sin wt]
e (max) = eo, when sin wt = 1 or angle wt = 90 degree
----------------------
eo = BANw
w = eo/BAN = 2.10*10-2/(0.0769*0.0155*0.0155*160)
w = 7.188236 rad/sec